OpenStudy (anonymous):

solving equations using factoring 3x^2 – 75 = 0

OpenStudy (anonymous):

factor out the 3 first and write \(3(x^2-25)=0\) then factor \(x^2-25\)

jimthompson5910 (jim_thompson5910):

use the rule a^2 - b^2 = (a-b)(a+b) to factor x^2-25